1% error with the method of dividing the root of the equation in the range [1,2]...

f(x) = x^{3}+4x^{2}-10  1% error with the method of dividing the root of the equation in the range [1,2] by two Solve it?
Answer #1

e feay = x 422 10 - 0 = S-10 C-5 f0 = U 920 (2) = 2 F4, 2 - - 0 = 8+ /6 - 10 = 24-10 = 14 fy, f(2) - - 5x19 = -7ㅇ ㅇ, so root